OpenStudy (anonymous):

to prepare 300.0 mL of 0.2 M NaCl, how many grams of NaCl should be weighed out?

OpenStudy (aaronq):

use the molarity formula: \(M=\dfrac{n_{solute}}{L_{solution}}\) you can solve for moles then convert to mass, or substitute in \(n=\dfrac{m}{M_{molar}}\) \(M=\dfrac{\dfrac{m}{M_{molar}}}{L_{solution}}=\dfrac{m}{M_{molar}*L_{solution}}\) Rearranges to: \(m=M*M_{molar}*L_{solution}\) Where: \(M_{molar}\)=molar mass m=mass L=liters of solution n=moles M=molarity

OpenStudy (somy):

@aaronq awesome work as usual!!!!!!!!

OpenStudy (anonymous):

thank you!!!